If a certain silver wire has a resistance of 6.20 2 at 40.0°C, what resistance will it have at 55.0°C?
Answer #1

The resistance is, R =R(1+aſt -t)) =6.20(1+(0.0038)(55-40)) = 6.55340 =6.550
Answer #2

SOLUTION :


Temperature coefficient of resistance of silver “alpha” is 0.0038 per ºC .


Resistance increases with increase in temperature and it is given by the following formula :


R = Ro ( 1 + alpha * temperature difference)

=> R55 = 6.20 ( 1 + 0.0038 * (55 - 40))

=> R55 = 6.5534 Ω  = 6.6 Ω approx.  (ANSWER).

answered by: Tulsiram Garg